2 Calculating a Pareto optimal allocation Suppose the representative household has the following utility function: U (C,) InC +0.5ln l where C is consumption and 1 is leisure. The households time constraint is I+N-1 where Ns is the representative households labour supply. Further assume that the production function is Cobb-Douglas 0.5 0.5 where 2-1 and K = 1 2.1 Assuming that the government spending is G = 0, use the Social Planners problem to solve for Pareto optimal numerical values C.r.N*.Y (8 points) 2.2 Find the numerical values of the real wage rate and of the representa- tive firms profit π* in a competitive equilibrium implementing the Social Planners (Pareto) optimal solution (4 points) 2.3 Draw a carefully labeled chart showing the impact of an increase in G on the Pareto Optimal allocation. Your chart does not have to be drawn exactly to scale. (7 points) 2.4 How would the optimal values C *,N. Y*,w*, π* dange (in what di- rection) in case if government spending increases (6 points)?
Answer #1

The social planner has to maximise the consumer's utility. So the social planner solves:

max   u(C,N)
subject to. C = f(l)
and   l + N = 1

In order to solve the planning problem, we use Lagrange multipliers. Let phi_{t} denote the Lagrange multiplier on the time constraint and let phi_{c} denote the Lagrange multiplier on the consumption constraint.

Since there are two constraints, the Lagrangean for this problem is given by:
L(C,N,l,phi_{c},phi_{t}) = u(C,N) + phi_{c}(f(l) - c) + phi_{t}(1 - l - N)

For our problem, the equations are given as:
max In(C) + 0.51n(1) = In(C)0.5n- N)
subject to. C = (N)12 = 1-1
  l + N = 1

The first order conditions for this problem are given by:
[C] : rac{partial u}{partial C} = phi_{c}^{*} = 1/C
[N] : àu = phi_{t}^{*} = 0.5

[l] :  phi_{c}^{*}imes (1/2)(1 - l^{*})^{-1/2} = phi_{t}^{*}
[phi_{c}] : C^{*} = (1 - l^{*})^{1/2}
[phi_{t}] : l^{*} + N^{*} = 1

The superscript * indicates that the quantities are the optimal quantities of the planning problem.
Thus, solving, we get optimal quantities as
I* =1/2
N^{*} = 1/2
C^{*} = (1/2)^{0.5}
The value of  phi_{t}^{*}/ phi_{c}^{*} gives us a value precisely equal to the equilibrium real wage.

When government spending exists, the equation of C = f(l) gets an additional factor, and becomes
C = f(l) - G
It affects the equilibrium optimal conditions as shown in the graph (the value of C decreases by Gz and the value of l decreases by (G/z):